21.  The quadratic equation whose roots are -1/2 and 2/3 is?


1
Expert's answer
2021-02-28T14:26:06-0500

a(x+1/2)(x-2/3)=ax2-ax/6-a/3

Answer: ax2-ax/6-a/3=0 , a "\mathrlap{\,/}{=}" 0. For example: if a=1, then x2-x/6-1/3=0